OpenStudy (laddiusmaximus):

pi^(5)(3x-1) calculate the derivative.

OpenStudy (laddiusmaximus):

I know the product rule is used, but Im getting lost.

OpenStudy (laddiusmaximus):

I thought the derivative of pi^5 is 5pi^4?

OpenStudy (kinggeorge):

Since pi is a constant, you can multiply it out, so you're taking the derivative of \[\pi^5 3x -\pi^5\]Which is just \(3\pi^5\) by the power rule.

OpenStudy (kinggeorge):

Since \(\pi\) is a constant, \({d\over dx}\;\;\pi^5 = 0\)

OpenStudy (laddiusmaximus):

yeah thats not right

OpenStudy (kinggeorge):

Are you deriving with respect to x or pi?

OpenStudy (kinggeorge):

In other words, is pi a variable, or are you using it as the constant 3.14159...?

Directrix (directrix):

y' = 3 π ^5